Solve for x: 3 + cos 2x = 7/2; only include answers on [0, 2π]
Answer by lwsshak3(11628)

Solve for x: 3 + cos 2x = 7/2; only include answers on [0, 2π]
3+cos2x=7/2
cos2x=3.5-3
cos2x=0.5
2x=π/3, 5π/3
x=π/6, 5π/6